# ViSTA: Visual Storytelling using Multi-modal Adapters for Text-to-Image Diffusion Models

### v2

@@ 3.5. Text-Image Alignment Evaluation with TIFA
Figure 4. Qualitative results: StorySalon. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. We include ground truth images from the dataset as a reference for evaluating visual coherence and accuracy. The same reference image is used across IP-Adapter, StoryGen, and ViSTA. The first story prompt corresponds to the reference image, while story prompts 2 to 7 are used to generate the subsequent frames. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
4. Experiments
4.1. Experimental Setup
Dataset.
We conduct our experiments on two dataset, StorySalon (Liu et al., 2024) and FlintStonesSV (Maharana and Bansal, 2021). The StorySalon dataset (Liu et al., 2024) is the largest and most recent benchmark for visual storytelling. It contains 159,778 animation-style images across 446 character categories. Designed to support long-range story generation, each story consists of an average of 14 frames, and each corresponding text prompt contains an average of 106 words. The dataset is constructed from E-books and YouTube videos retrieved using keyword-based queries. To generate the narrative-aligned text descriptions for each image, TextBind (Li et al., 2024) is applied to produce captions conditioned on both the image and the surrounding narrative text.
Due to a portion of the YouTube videos becoming unavailable, we conduct all experiments on the E-books partition, which remains complete and high-quality. This split includes 8,635 training stories with 118,892 frames, and 451 test stories with 6,026 frames.
The FlintstonesSV dataset (Maharana and Bansal, 2021) serves as an additional benchmarking dataset in our experiments. It comprises 25,184 densely annotated, animation-style video clips sourced from the animated series “The Flintstones.” FlintstonesSV selects a single representative frame from each clip and groups frames from consecutive clips into coherent stories of length five. The dataset consists of seven recurring characters, offering a robust setting for evaluating character and scene consistency in visual storytelling tasks. FlintstonesSV is split into 20,132 training, 2,071 validation, and 2,309 test stories.
@@ 4.1. Experimental Setup
Model CLIP-T $\uparrow$ TIFA $\uparrow$ CLIP-I $\uparrow$ FID $\downarrow$
SDXL-Prompt 0.319 0.782 0.697 63.66
IP-Adapter 0.255 0.521 0.725 42.66
StoryGen (Liu et al., 2024) 0.313 0.750 0.710 47.28
ViSTA (ours) 0.316 0.765 0.715 46.45
@@
Table 1. Quantitative results. We evaluate text-image alignment and image-image consistency on the E-Book partition of the StorySalon test set. The SDXL-Prompt baseline, which generates images using only the current text prompt with an animation prefix, achieves the highest text-image alignment scores (CLIP-T and TIFA). However, it suffers from the lowest image-image similarity (CLIP-I) and fidelity (FID) due to the lack of visual conditioning–i.e., SDXL model only takes the current prompt as input. In contrast, IP-Adapter, which conditions image generation solely on history images along with the current prompt, significantly improves image-image similarity and FID, as it directly references visual appearance from previous frames. However, it performs worse on text-image alignment because it lacks access to the full narrative prompt and relies only on visual cues without detailed textual guidance. These results highlight the importance of conditioning on both history text and image features to achieve a balance between visual consistency and narrative relevance. Our proposed ViSTA model, which leverages both modalities, outperforms the state-of-the-art StoryGen (Liu et al., 2024) on all evaluation metrics. Blue bold numbers indicate the best scores, red numbers indicate the worst, and underlined numbers represent state-of-the-art result for storytelling.
Evaluation Metrics.
To evaluate text-image alignment, we utilize CLIP text-image similarity (CLIP-T (Radford et al., 2021)) and TIFA (Hu et al., 2023) score. Following the literature (Feng et al., 2023; Pan et al., 2024; Liu et al., 2024), we use CLIP-T to evaluate high-level semantic text-image similarity. We also propose to use TIFA to assess the how well the generated images align with the narrative story text. Following (Hu et al., 2023), we use GPT-3.5 (Brown et al., 2020) to generate question-answer pairs and use UnifiedQA (Khashabi et al., 2020) to evaluate generated images. [^1] https://github.com/Yushi-Hu/tifa
To evaluate the quality of the generated images with respect to the ground-truth, we use Frechet Inception Distance (FID (Heusel et al., 2017)) score and CLIP image-image similarity (CLIP-I (Radford et al., 2021)). FID is used to evaluate the quality of generated images by comparing the distributions of generated images and ground-truth images, quantifying the faithfulness and diverseness of generated images. CLIP-I measures the similarity between the generated and ground-truth images.
Baselines.
For StorySalon dataset, we compare our method with the following: (1) SDXL-Prompt: [^2] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 We use the pretrained Stable Diffusion XL model (Podell et al., 2024) given the current text as prompt with a “A cartoon style image” prefix, without using any history image or text. (2) IP-Adapter (Ye et al., 2023): We use the IP-Adapter for SDXL [^3] https://huggingface.co/h94/IP-Adapter/tree/main/sdxl_models to generate the image given the current text as its prompt, and all the previous generated history images as its reference images. (3) State-of-the-art StoryGen (Liu et al., 2024): We run the inference using the public checkpoint of the StoryGen model. [^4] https://huggingface.co/haoningwu/StoryGen/tree/main/checkpoint_StorySalon For FlintStonesSV datset, we compare our method with SDXL-Prompt, IP-Adapter and StoryGPT-V (Shen and Elhoseiny, 2025).
Implementation Details.
Our multi-modal history fusion model consists of $d=4$ blocks, each block contains a cross-attention layer and a FeedForward Network. The hidden dimension of the fusion model is 1024. We use OpenCLIP-ViT-H (Cherti et al., 2023) as the image and text encoders. Our history adapter is built on the Stable Diffusion XL. We train the model end-to-end for 80,000 steps with a batch size of 4. We use AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of 1e-4. During inference, we adopt a DDIM sampler (Song et al., 2021) with 50 inference steps, and the guidance scale is 5. All experiments are conducted on 4 NVIDIA L40S GPUs. We set $\lambda=0.5$ to balance image quality and consistency. The larger $\lambda$ improves consistency but reduces quality, while the smaller $\lambda$ improves quality at the cost of consistency. The value $0.5$ offers the best trade-off in our experiments.
4.2. Results
@@
We evaluate our proposed method on two different dataset with different settings, StorySalon (Liu et al., 2024) which does not contain recurring characters, while FlintStonesSV (Maharana and Bansal, 2021) has seven recurring characters. For both dataset, the first text-image pair is given as an initial reference. Afterwards, the generated frames serve as history for further generations. Table 1 presents the quantitative results on the E-book partition of the StorySalon test dataset, which contains 451 stories and 14 frames per story on average. As shown, SDXL-Prompt achieves the highest scores on CLIP-T and TIFA, demonstrating strong alignment with the input text prompt. However, without access to the history image or text, it performs poorly on CLIP-I and FID, indicating low visual consistency across frames. In contrast, IP-Adapter, which is conditioned solely on history images, achieves significantly better image-image similarity and fidelity but at the cost of degraded text-image alignment, due to the absence of detailed textual guidance of the history. The state-of-the-art method, StoryGen, provides a more balanced performance, but it still struggles to fully leverage both text and image history.
@@
Our proposed model, ViSTA, which conditions history text and image, outperforms StoryGen, a state-of-the-art model, across all metrics, demonstrating its effectiveness in preserving both text alignment and image consistency. Moreover, ViSTA employs a lightweight adapter without modifying the diffusion model architecture or requiring full fine-tuning, further highlighting its efficiency and practical applicability in generating high-quality story images.
@@
Figure 4 shows the qualitative results of ViSTA, baselines, and state-of-the-art methods. The same reference image is initially used across IP-Adapter, StoryGen, and ViSTA. The first story prompt is the corresponding caption for the reference image and is used for both StoryGen and ViSTA. We also include the ground truth images from the StorySalon dataset, which serve as the reference for evaluating the coherence and accuracy of the generated story images. These real images illustrate how the same character and style evolve naturally throughout the story, providing a target for visual and narrative consistency. We observe that although SDXL-Prompt generates high-quality and well-aligned images, it struggles to maintain character consistency across frames. In contrast, IP-Adapter ensures consistent character appearances but lacks strong text-image alignment, often failing to accurately reflect the given prompts. Compared to the state-of-the-art StoryGen, ViSTA achieves superior consistency in both character identity and visual style, effectively improving text alignment and consistent storytelling throughout the generated image sequence.
@@
Table 2 presents quantitative results on the FlintstonesSV test dataset. Our observations on FlintstonesSV align closely with previous findings on SDXL-Prompt and IP-Adapter, demonstrating ViSTA’s effectiveness in preserving both text alignment and image consistency. Compared to the other state-of-the-art methods, StoryGPT-V (Shen and Elhoseiny, 2025) , ViSTA achieves superior results, indicating notably improved quality in terms of image realism and distribution similarity. More results are presented in the appendix.
@@
Table 2. Quantitative results: FlintStonesSV. StoryGPT-V numbers are reported in (Shen and Elhoseiny, 2025). Results are consistent with prior SDXL-Prompt and IP-Adapter observations. ViSTA outperforms SOTA StoryGPT-V (Shen and Elhoseiny, 2025) on FID on the recurring characters setting, indicating improved image quality.
5. Conclusion
@@
In this paper, we present ViSTA, a novel approach for visual storytelling that enhances character consistency and text-image alignment in text-to-image diffusion models. To effectively utilize history context, we introduce a multi-modal history fusion model that extracts and integrates relevant features from past text-image pairs, enabling coherent story generation. Additionally, we propose a lightweight history adapter that conditions the diffusion model on history fusion features without modifying its architecture or requiring full fine-tuning, ensuring efficiency and adaptability. To further improve the generation process, we introduce a salient history selection strategy that dynamically identifies the most informative history text-image pair at each step, enhancing the quality of conditioning. Furthermore, we are the first to employ TIFA as a metric for evaluating text-image alignment in visual storytelling, providing a more targeted and interpretable assessment of generated images. Extensive experiments on two datasets demonstrate that ViSTA outperforms existing state-of-the-art methods, achieving superior character consistency and text-image alignment.
Appendix A Additional Results
A.1. StorySalon Dataset
@@
Additional qualitative results from the StorySalon dataset are presented in Figures 7, 8, 9, and 10. Consistent patterns are observed across different stories, including those featuring both human characters and animals, demonstrating that ViSTA achieves superior character and style consistency compared to other state-of-the-art baseline methods.
A.2. FlintStonesSV Dataset
@@
Figures 5 and 6 illustrate examples generated by ViSTA from the FlintstonesSV dataset. The results demonstrate ViSTA’s effectiveness in maintaining consistent character identity and coherent visual storytelling across the sequence of images.
Appendix B Limitations

### v3

@@ 3.5 Text-Image Alignment Evaluation with TIFA
Figure 4: Qualitative results: StorySalon. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. We include ground truth images from the dataset as an evaluation for visual coherence and accuracy. The same reference image is used across all methods. The first story prompt corresponds to the reference image, while story prompts 2 to 7 are used to generate the subsequent frames. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character across all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
4 Experiments
We evaluate our proposed method on two different dataset with different settings, StorySalon [13] which does not contain recurring characters, while FlintStonesSV [17] has seven recurring characters. For both datasets, the first text-image pair is given as an initial reference. Afterwards, the generated frames serve as history for further generations. Experimental setup, including dataset descriptions, evaluation metrics, baselines, implementation details, as well as the code can be found in the supplementary material.
@@ 4 Experiments
Model CLIP-T $\uparrow$ TIFA $\uparrow$ CLIP-I $\uparrow$ FID $\downarrow$ Inference Time (s)
SDXL-Prompt 0.319 0.782 0.697 63.66 5.83
IP-Adapter 0.255 0.521 0.725 42.66 6.41
StoryGen [13] 0.313 0.750 0.710 47.28 8.23
ViSTA (ours) 0.316 0.765 0.715 46.45 6.33
@@
Table 1: Quantitative results. We evaluate text-image alignment and image-image consistency on the E-Book partition of the StorySalon test set. The SDXL-Prompt baseline, which generates images using only the current text prompt with an animation prefix, achieves the highest text-image alignment scores (CLIP-T and TIFA). However, it suffers from the lowest image-image similarity (CLIP-I) and fidelity (FID) due to the lack of visual conditioning–i.e., SDXL model only takes the current prompt as input. In contrast, IP-Adapter, which conditions image generation solely on history images along with the current prompt, significantly improves image-image similarity and FID, as it directly references visual appearance from previous frames. However, it performs worse on text-image alignment because it lacks access to the full narrative prompt and relies only on visual cues without detailed textual guidance. These results highlight the importance of conditioning on both history text and image features to achieve a balance between visual consistency and narrative relevance. Our proposed ViSTA model, which leverages both modalities, outperforms the state-of-the-art StoryGen [13] on all evaluation metrics. Blue bold numbers indicate the best scores, red numbers indicate the worst, and underlined numbers represent state-of-the-art result for storytelling.
4.1 Results
@@
Table 1 presents the quantitative results on the E-book partition of the StorySalon test dataset. As shown, SDXL-Prompt achieves the highest scores on CLIP-T and TIFA, demonstrating strong alignment with the input text prompt. However, without access to the history image or text, SDXL-Prompt performs poorly on CLIP-I and FID, indicating low visual consistency across frames. In contrast, IP-Adapter, which is conditioned solely on history images, achieves significantly better image-image similarity and fidelity but at the cost of degraded text-image alignment, due to the absence of detailed textual guidance of the history. The state-of-the-art method, StoryGen, provides a more balanced performance, but it still struggles to fully leverage both text and image history.
@@
Our proposed model, ViSTA, which conditions history text and image, outperforms StoryGen, a state-of-the-art model, across all metrics, demonstrating the effectiveness in preserving both text alignment and image consistency. Moreover, ViSTA employs a lightweight adapter without modifying the diffusion model architecture or requiring full fine-tuning, further highlighting its efficiency and practical applicability in generating high-quality story images.
We also compared the inference time (in seconds) for each method under the same hardware setting (one L40S GPU). Note that StoryGen is based on SD v1.5, which uses a smaller backbone than SDXL but incorporates an auto-regressive generation module that has a significant 7.3x increase on inference time. In contrast, our method is built on SDXL and introduces only a lightweight history adapter, resulting in a minimal increase in runtime–just 1.1x (0.5 seconds) more than vanilla SDXL. Compared to IP-Adapter, ViSTA achieves comparable efficiency while delivering improved character consistency and narrative alignment.
@@
Figure 4 shows the qualitative results of ViSTA, baselines, and state-of-the-art methods. The same reference image is initially used across IP-Adapter, StoryGen, and ViSTA. The first story prompt is the corresponding caption for the reference image and is used for both StoryGen and ViSTA. We also include the ground truth images from the StorySalon dataset, which serve as the reference for evaluating the coherence and accuracy of the generated story images. These real images illustrate how the same character and style evolve naturally throughout the story, providing a target for visual and narrative consistency. We observe that although SDXL-Prompt generates high-quality and well-aligned images, it struggles to maintain character consistency across frames. In contrast, IP-Adapter ensures consistent character appearances but lacks strong text-image alignment, often failing to accurately reflect the given prompts. Compared to the state-of-the-art StoryGen, ViSTA achieves superior consistency in both character identity and visual style, effectively improving text alignment and consistency throughout the generated image sequence.
@@
Table 2 presents quantitative results on the FlintstonesSV test dataset. Our observations on FlintstonesSV align closely with previous findings on SDXL-Prompt and IP-Adapter, demonstrating ViSTA’s effectiveness in preserving both text alignment and image consistency. Compared to the other state-of-the-art methods, StoryGPT-V [27] , ViSTA achieves superior results, indicating notably improved quality in terms of image realism and distribution similarity. More results are presented in the appendix.
@@
Table 2: Quantitative results: FlintStonesSV. StoryGPT-V numbers are reported in [27]. Results are consistent with prior SDXL-Prompt and IP-Adapter observations. ViSTA outperforms SOTA StoryGPT-V [27] on FID on the recurring characters setting, indicating improved image quality.
4.2 Human Evaluation
We conducted a pilot study to rank the story generation methods on three main criteria: (1) Textual Alignment: how well does each frame follow its corresponding prompt (2) Image Alignment: how well does each frame follow the overall style of the reference image, and (3) Consistency: how well is the frame-to-frame consistency (same characters/objects). We asked five non-author graduate students to rank five examples per person. Table 3 shows the elo ranking. The initial elo score was 128, and the tuning factor is 4. It is also worth noting that our method is the top 1 ranked method 69.23%, 57.69%, and 50.00% of the time when compared with the other 3 methods in Textual Alignment, Image Alignment, and Consistency, respectively. More detail can be found in the supplementary material.
Model Textual Align. Image Align. Consistency
SDXL-Prompt 174.78 128.51 117.77
IP-Adapter 8.62 81.21 178.05
StoryGen 107.71 85.00 26.03
ViSTA (ours) 220.88 217.28 190.15
Table 3: User Study. Elo ranking for all methods across Textual Alignment, Image Alignment, and Consistency metrics.
5 Conclusion
@@
In this paper, we present ViSTA, a novel approach for visual storytelling that enhances character consistency and text-image alignment in text-to-image diffusion models. To effectively utilize history context, we introduce a multi-modal history fusion model that extracts and integrates relevant features from past text-image pairs, enabling coherent story generation. Additionally, we propose a lightweight history adapter that conditions the diffusion model on history fusion features without modifying its architecture or requiring full fine-tuning, ensuring efficiency and adaptability. To further improve the generation process, we introduce a salient history selection strategy that dynamically identifies the most informative history text-image pair at each step, enhancing the quality of conditioning. Furthermore, we are the first to employ TIFA as a metric for evaluating text-image alignment in visual storytelling, providing a more targeted and interpretable assessment of generated images. Extensive experiments on two datasets demonstrate that ViSTA outperforms state-of-the-art methods, achieving superior character consistency and text-image alignment.
Appendix A Experimental Setup
A.1 Datasets
We conduct our experiments on two dataset, StorySalon [13] and FlintStonesSV [17]. The StorySalon dataset [13] is the largest and most recent benchmark for visual storytelling. It contains 159,778 animation-style images across 446 character categories. Designed to support long-range story generation, each story consists of an average of 14 frames, and each corresponding text prompt contains an average of 106 words. The dataset is constructed from E-books and YouTube videos retrieved using keyword-based queries. To generate the narrative-aligned text descriptions for each image, TextBind [12] is applied to produce captions conditioned on both the image and the surrounding narrative text. Due to a portion of the YouTube videos becoming unavailable, we conduct all experiments on the E-books partition, which remains complete and high-quality. It includes 8,635 training stories with 118,892 frames, and 451 test stories with 6,026 frames.
The FlintstonesSV dataset [17] serves as an additional benchmarking dataset in our experiments. It comprises 25,184 densely annotated, animation-style video clips sourced from the animated series “The Flintstones.” FlintstonesSV selects a single representative frame from each clip and groups frames from consecutive clips into coherent stories of length five. The dataset consists of seven recurring characters, offering a robust setting for evaluating character and scene consistency in visual storytelling tasks. FlintstonesSV is split into 20,132 training, 2,071 validation, and 2,309 test stories.
A.2 Evaluation Metrics
To evaluate text-image alignment, we utilize CLIP text-image similarity (CLIP-T [21]) and TIFA [9] score. Following the literature [4, 19, 13], we use CLIP-T to evaluate high-level semantic text-image similarity. We also propose to use TIFA to assess the how well the generated images align with the narrative story text. Following [9], we use GPT-3.5 [2] to generate question-answer pairs and use UnifiedQA [11] to evaluate generated images.  [^1] https://github.com/Yushi-Hu/tifa
To evaluate the quality of the generated images with respect to the ground-truth, we use Frechet Inception Distance (FID [7]) score and CLIP image-image similarity (CLIP-I [21]). FID is used to evaluate the quality of generated images by comparing the distributions of generated images and ground-truth images, quantifying the faithfulness and diverseness of generated images. CLIP-I measures the similarity between the generated and ground-truth images.
A.3 Baselines
For StorySalon dataset, we compare our method with the following: (1) SDXL-Prompt:  [^2] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 We use the pretrained Stable Diffusion XL model [20] given the current text as prompt with a “A cartoon style image” prefix, without using any history image or text. (2) IP-Adapter [36]: We use the IP-Adapter for SDXL  [^3] https://huggingface.co/h94/IP-Adapter/tree/main/sdxl_models to generate the image given the current text as its prompt, and all the previous generated history images as its reference images. (3) State-of-the-art StoryGen [13]: We run the inference using the public checkpoint of the StoryGen model.  [^4] https://huggingface.co/haoningwu/StoryGen/tree/main/checkpoint_StorySalon For FlintStonesSV datset, we compare our method with SDXL-Prompt, IP-Adapter and StoryGPT-V [27].
While a wide range of methods exist for visual storytelling, many rely on assumptions or constraints that are incompatible with our broader setting.
Specifically, we compare against StoryGen on the StorySalon dataset and StoryGPT-V on the FlintstonesSV dataset, which are, to the best of our knowledge, the only methods that operate under our broader setting.
Other methods were not included for the following reasons: (1) StoryGPT-V leverages predefined recurring characters specific to the FlintstonesSV dataset, a feature not available in the StorySalon dataset. This makes it inapplicable for evaluation on StorySalon. (2) We also excluded certain training-free methods (e.g., StoryDiffusion) because they are designed for short, “character + activity” format prompts and rely on repeated character prompt for consistency. We did experiments on StorySalon dataset with StoryDiffusion and OnePromptOneStory methods in our setting but yielded poor results. We believe that such comparisons would not be fair or meaningful.
Therefore, to our best knowledge, we chose the baselines that do not make such assumptions and are therefore applicable to the datasets we compare against.
A.4 Implementation Details
Our multi-modal history fusion model consists of $d=4$ blocks, each block contains a cross-attention layer and a FeedForward Network. The hidden dimension of the fusion model is 1024. We use OpenCLIP-ViT-H [3] as the image and text encoders. Our history adapter is built on the Stable Diffusion XL. We train the model end-to-end for 80,000 steps with a batch size of 4. We use AdamW optimizer [15] with a learning rate of 1e-4. During inference, we adopt a DDIM sampler [29] with 50 inference steps, and the guidance scale is 5. All experiments are conducted on 4 NVIDIA L40S GPUs. We set $\lambda=0.5$ to balance image quality and consistency. The larger $\lambda$ improves consistency but reduces quality, while the smaller $\lambda$ improves quality at the cost of consistency. The value $0.5$ offers the best trade-off in our experiments.
Appendix B Additional Results
B.1 StorySalon Dataset
@@
Additional qualitative results from the StorySalon dataset are presented in Figures 9, 10, 11, and 12. Consistent patterns are observed across different stories, including those featuring both human characters and animals, demonstrating that ViSTA achieves superior character and style consistency compared to other state-of-the-art baseline methods.
B.2 FlintStonesSV Dataset
@@
Figures 5 and 6 illustrate examples generated by ViSTA from the FlintstonesSV dataset. The results demonstrate ViSTA’s effectiveness in maintaining consistent character identity and coherent visual storytelling.
Figure 5: ViSTA sample on FlintStonesSV test set.
Figure 6: ViSTA sample on FlintStonesSV test set.
Appendix C Human Evaluation
To evaluate the quality of the generated story sequences, we conducted a pilot human study with five graduate student participants. Each participant was asked to assess the outputs from four different story generation methods.
The evaluation began with a set of detailed instructions provided through the survey interface, as shown below. These instructions outlined the evaluation procedure and defined the three core criteria: textual alignment, image alignment, and consistency.
Story Generation Methods Assessment In this survey you will compare 4 different story generation methods that use state-of-the-art stable diffusion models across 5 examples. All methods start with a single prompt with its corresponding image as reference. After that, they generate a single image for each subsequent story prompt. For each example you will see the reference prompt-image pair and the subsequent story prompts, in addition to the original story book for reference on how the story looks like in a real-life example. You will be asked to compare the generated story based on 3 factors/metrics: 1. Textual Alignment: how well does each frame follow its corresponding prompt 2. Image Alignment: how well does each frame follow the overall style of the reference image 3. Consistency: how well is the frame-to-frame consistency (same characters/objects) You will rank each metric separately for all methods giving 1 to the best method and 4 to the worst. Below are some examples of good/bad generations based on each metric.
Figure 7 presents the detailed explanations of each evaluation criterion, along with visual examples illustrating both high-quality and low-quality generations. These examples were included to calibrate participants’ understanding and ensure consistent evaluation across annotators.
Figure 8 illustrates the evaluation interface and procedure. For each of the five examples, participants were presented with a reference prompt and image, as well as the original story images. They then reviewed the image sequences generated by all four methods, shown in randomized order. For each criterion, participants independently ranked the four methods from 1 (best) to 4 (worst).
Each participant evaluated five examples, resulting in a total of 25 sets of rankings per metric across the study.
Figure 7: Examples illustrating good and poor outputs for each of the three evaluation metrics: textual alignment, image alignment, and consistency. These examples were shown to participants to calibrate their understanding of the evaluation criteria.
Figure 8: Interface for the human evaluation study. Each example included a reference image and prompt, original story content, and the outputs from four different generation methods (only showing one method for brevity). Participants ranked the methods for each metric independently (only showing one metric for brevity).
Appendix D Limitations